Joanne Germano works in an accounts payable department of a major retailer. She has attempted to convince her boss to take the discount on the 2 divided by 15 net 30 credit terms most suppliers​ offer, but her boss argues that giving up the 2​% discount is less costly than a​ short-term loan at 11​%. Prove to whoever is wrong that the other is correct.  

​(​Note: Assume a​ 365-day year.)
Answer #1

1. The terms 2/15 net 30 means that if payment is done within 15 days tehn a discoune of 2% will be offered otherwise payment has to be done in 30 days.

forgoing discount means paying 2% extra by paying 15 days later

Hence One period consists of 15 days.

Periodic rate = 2%

Effective Annual Rate = (1+i)n - 1

Here, i = periodic rate

n = number of periods in a year

Effective cost = ( 1+0.02)(365÷15) - 1 = 0.6191 = 61.91%

Effective cost of borrowing = 11%

Hence borrowing and taking the discount is less costly to the company.

Joanne Germano is correct in saying to her boss that borrowing and taking a discount is better option for the company.
Answer #2
I think its 45×3% then we add to the short term loan
source: Corporate finance notes
answered by: tina